OpenStudy (anonymous):

i need help to verify (tan^2x)/(1-cosx) = secx+sec^2x

OpenStudy (anonymous):

the trig identity for tangent ^2 is sec^-1

OpenStudy (anonymous):

sec^2-1

OpenStudy (anonymous):

okay yea that makes sense so far.

OpenStudy (zale101):

I'm not sure if the trig identity for tangent ^2 is sec^2x-1..would it be \(tan^{2}x=\Large \frac{sin^2x}{cos^2x}\)?

OpenStudy (anonymous):

well the pythagorean identity for tangent^2 is sec^2x-1 at least that is what my teacher gave me idk if i can do it the other way. but that might make it easier.

OpenStudy (zale101):

brb, i'll try to ask someone

OpenStudy (anonymous):

okie dokes

OpenStudy (zale101):

Ok, it's correct

OpenStudy (zale101):

I apologize for taking too long

OpenStudy (anonymous):

okay cool its fine what would i do from there.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\tan^2(x)}{1-\cos(x)} = \sec(x) + \sec^2(x)\] \[\Large \frac{\sec^2(x)-1}{1-\cos(x)} = \sec(x) + \sec^2(x)\] \[\Large \frac{\frac{1}{\cos^2(x)}-1}{1-\cos(x)} = \sec(x) + \sec^2(x)\] \[\Large \frac{\cos^2(x)}{\cos^2(x)}*\frac{\frac{1}{\cos^2(x)}-1}{1-\cos(x)} = \sec(x) + \sec^2(x)\] \[\Large \frac{\cos^2(x)(\frac{1}{\cos^2(x)}-1)}{\cos^2(x)(1-\cos(x))} = \sec(x) + \sec^2(x)\] \[\Large \frac{1-\cos^2(x)}{\cos^2(x)(1-\cos(x))} = \sec(x) + \sec^2(x)\] \[\Large \frac{(1-\cos(x))(1+\cos(x))}{\cos^2(x)(1-\cos(x))} = \sec(x) + \sec^2(x)\] Hopefully you see what to do from here

OpenStudy (anonymous):

i kind of do but im not to good at this. could you help me finish i know you would use reciprocal identities but..